a rectangle has a lenght which is 5cm less than twice its breadth. if the length is decreased by 5 cm and breadth is increased by 2 cm.the perimeter of the resulting rectangle will be 74.find the length and breadth
Answer by Cromlix(4381)

Breadth = x
Length = 2x - 5
Then the length is decreased by 5
= 2x - 10
The breadth is increased by 2
= x + 2
Perimeter = 2*breadth + 2*length
74 = 2*(x + 2)+ 2*(2x - 10)
74 = 2x + 4 + 4x - 20
74 = 6x - 16
6x = 74 + 16
6x = 90
x = 15 cm
Breadth = 15cm
Length = 25cm
Hope this helps.
:-)